Determine the solubility of the following in mol/L at 20°C.

a) NH4Cl ______in cold water and ______hot water b)Na2CO3 ______in cold water and_____ hot water. Kindly show calculations
1
Expert's answer
2020-02-25T04:54:08-0500

a) s20=372.0 g/L (20°C)

s100=740.8 g/L (100 °C)

M(NH4Cl)=53 g/mol

S20=s20/M=372/53=7 mol/L

S100=s100/M=740.8/53=14 mol/L

b) s20=218.0 g/L (20 °C)

s100=436.0 g/L (100 °C)

M(Na2CO3)=106 g/mol

S20=s20/M=218/106=2.06 mol/L

S100=s100/M=436/106=4.11 mol/L